Solve the given equation and identify it as a conditional equation, an inconsistent equation, or an identity.
(8 – 8)(7 – p) = 0

Answer by mangopeeler07(462)

(8 – 8)(7 – p) = 0
This is an identity. Check it out:

(8-8)=0
(7-p)=(7-p)

Plug in these values into (8-8)(7-p)=0
(0)(7-p)=0

This is saying "zero times a quantity equals zero". This is always true, no matter what the quantity is, which makes this an identity.